# Colorful Pinball: Density-Weighted Quantile Regression for Conditional Guarantee of Conformal Prediction

### v3

@@ Abstract
While conformal prediction provides robust marginal coverage guarantees, achieving reliable conditional coverage for specific inputs remains challenging. Although exact distribution-free conditional coverage is impossible with finite samples, recent work has focused on improving the conditional coverage of standard conformal procedures. Distinct from approaches that target relaxed notions of conditional coverage, we directly minimize the mean squared error of conditional coverage by refining the quantile regression components that underpin many conformal methods. Leveraging a Taylor expansion, we derive a sharp surrogate objective for quantile regression: a density-weighted pinball loss, where the weights are given by the conditional density of the conformity score evaluated at the true quantile. We propose a three-headed quantile network that estimates these weights via finite differences using auxiliary quantile levels at $1-\alpha\pm\delta$, subsequently fine-tuning the central quantile by optimizing the weighted loss. We provide a theoretical analysis with exact non-asymptotic guarantees characterizing the resulting excess risk. Extensive experiments on diverse high-dimensional real-world datasets demonstrate remarkable improvements in conditional coverage performance.
@@ 1 Introduction
As machine learning systems are increasingly deployed in high-stakes environments, the demand for reliability extends beyond predictive accuracy to rigorous uncertainty quantification (UQ). In safety-critical domains, a model’s ability to provide appropriate predictive intervals is often as valuable as accurate point predictions. Among existing UQ frameworks, Conformal Prediction (CP; Vovk et al., 2005, Shafer and Vovk, 2008) has emerged as a paradigm of choice due to its mathematically grounded guarantees. Unlike Bayesian methods or ensemble techniques that often rely on strong distributional assumptions or heavy computational overhead, CP offers a distribution-free, model-agnostic framework. It constructs prediction sets or intervals that provably contain the ground truth with a user-specified probability $1-\alpha$ (e.g., 90%) in finite samples, providing a layer of statistical trust that is indispensable for real-world deployment.
However, standard split conformal prediction only provides a marginal guarantee over the population and cannot guarantee conditional coverage for specific instances—precisely what practitioners require in high-stakes scenarios. Although hardness results for exact distribution-free conditional guarantees with finite samples are well-established Vovk [2012], Lei and Wasserman [2014], Foygel Barber et al. [2021], a growing body of works targets improving the conditional coverage of conformal procedures. To this end, many works explore to relax the conditional requirement Angelopoulos et al. [2024], e.g., group-conditional coverage Jung et al. [2023], Ding et al. [2023].
Instead, this paper targets controlling the exact conditional coverage rather than relaxed forms. A natural objective is the MSE of conditional coverage—termed Mean Squared Conditional Error (MSCE) in Kiyani et al. [2024]. We consider this metric since minimizing the MSCE provides high-probability control over conditional coverage deviations via Bernstein’s inequality. We build upon the connection between MSCE and the excess risk of the pinball loss established in recent literature Kiyani et al. [2024], Plassier et al. [2025a]. Specifically, these works show that the MSCE is upper-bounded by the excess risk of a quantile estimator with pinball loss, suggesting that improving the quantile estimator directly translates to better conditional coverage.
@@ 1.1 Related Works
We focus our discussion on the conditional coverage of conformal prediction. In the literature, there exists another branch of work concerning the training-conditional guarantee Park et al. [2020], Bian and Barber [2023], Duchi [2025], which studies the coverage probability conditional on the specific calibration set observed. To clarify, we are specifically concerned with the coverage guarantee conditional on the covariate of the test sample, $X_{\text{test}}$.
@@
Group-conditional coverage serves as a natural approximation to the exact conditional guarantee. Here, we specifically discuss the case where groups are formed by partitioning the covariate space. Classic conformal procedures operate with pre-defined groups Jung et al. [2023], where $h$ is a fixed discrete partition function. Beyond fixed groups, Kiyani et al. [2024] propose co-training the partition function with quantile regression within each group, providing a discrete approximation of conditional coverage. Though straightforward, the precision of this discrete approximation is limited, and the convergence rate of the resulting MSCE scales as slowly as $O(n^{-1/4})$. Adopting a different perspective that reformulates the conditional guarantee as moment equations, Gibbs et al. [2025a] extends the approximation to infinite, overlapping groups. They leverage regularized quantile regression as well as full conformal prediction Vovk et al. [2005]. However, Gibbs et al. [2025a] provides control only on a substitute metric—coverage under covariate shift—rather than the exact conditional coverage. Moreover, full conformal procedures incur a significant computational burden, limiting their general applicability.
@@
Beyond group-conditional guarantees, localized conformal prediction Bian and Barber [2023], Hore and Barber [2024] provides another avenue to approximate the conditional guarantee by applying kernel weighting during empirical quantile computation. As analyzed in Hore and Barber [2024], the guarantee provided by such localization corresponds to a random function $h$ that outputs a point sampled from a distribution centered at $X_{\text{test}}$, defined by the kernel. However, this strategy suffers inherently from the curse of dimensionality, particularly in the complex, high-dimensional scenarios that modern conformal prediction targets.
@@
Complementing the focus on grouping and quantile computation, a substantial body of literature strives to refine the nonconformity score function. Improving upon the classic residual score, Lei et al. [2018] propose scaled residuals using a learned conditional standard deviation $\sigma(x)$. Romano et al. [2019] propose Conformalized Quantile Regression (CQR), which directly conformalizes two learned quantiles rather than a mean prediction. They also highlighted the difficulty of learning $\sigma(x)$, particularly for overfitted neural networks. More recently, Xie et al. [2024] propose a boosting procedure to iteratively refine the score function. However, besides high computational costs, this method further requires access to training data, which is often infeasible given the prevalence of pre-trained black-box models. Similarly, CQR necessitates replacing the traditional conditional mean regression objective with quantile regression on the training set, thereby limiting its general applicability. While the aforementioned scores are tailored for one-dimensional labels, another branch of work investigates density-based scores Izbicki et al. [2022], Plassier et al. [2025b], Braun et al. [2025a], which are readily extensible to multi-dimensional labels and also flexible enough to capture heteroscedasticity. However, estimating the conditional density $f(Y\mid X)$ is generally a harder task than regression on the conditional mean $\mathbb{E}[Y\mid X]$. Furthermore, the calibration set used for density estimation is typically much smaller than the training set. This sample size mismatch limits the efficacy of density estimation; therefore, in this paper, we focus on the quantile regression of scores rather than density estimation.
@@
Most relevant to our work is the approach of performing quantile regression directly on conformity scores Gibbs et al. [2025a], Kiyani et al. [2024], Plassier et al. [2025a]. As aforementioned, the excess risk of the pinball loss constitutes an upper bound on the MSCE. Specifically, Plassier et al. [2025a] propose a general framework called Rectified Conformal Prediction (RCP), which transforms the original nonconformity score into a rectified version that tries to remove the covariate-dependent part in the score and thereby naturally improves conditional coverage of standard split conformal prediction with these rectified scores. However, naive quantile regression with the pinball loss faces intrinsic limitations, even for marginal coverage in high-dimensional settings Gibbs et al. [2025b]. In our context, we identify that naive quantile regression overlooks a critical heteroscedastic component. Our method aims to recover this missing component, yielding a significantly sharper approximation to the MSCE.
@@ 2 Preliminaries
We consider the standard setting where we observe data points $(X,Y)\in\mathcal{X}\times\mathcal{Y}$ drawn i.i.d. from an unknown joint distribution $P_{XY}$. In the split conformal prediction framework Papadopoulos et al. [2002], the available data is randomly partitioned into two disjoint subsets: a proper training set $\mathcal{D}_{\text{train}}$ and a calibration set $\mathcal{D}_{\text{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ of size $n$. A predictive model is first fitted on $\mathcal{D}_{\text{train}}$. We then define a nonconformity score function $S:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, which measures the discrepancy between the target $y$ and the model’s prediction at $x$ (e.g., the absolute residual $S(x,y)=|y-\hat{\mu}(x)|$). The scores are computed for all calibration points as $s_{i}=S(X_{i},Y_{i})$ for all $i\in\{1,\dots,n\}$. Given a user-specified miscoverage level $\alpha\in(0,1)$, we compute the conformal threshold $\hat{q}$ as the $\lceil(n+1)(1-\alpha)\rceil/n$-th empirical quantile of the calibration scores $\{s_{1},\dots,s_{n}\}$. For a new test input $X_{n+1}$, the prediction set is constructed as $\mathcal{C}(X_{n+1})=\{y\in\mathcal{Y}:S(X_{n+1},y)\leq\hat{q}\}$. This procedure satisfies the finite-sample marginal coverage guarantee:
@@
As has been widely noted, the failure pattern of standard split conformal prediction is under-coverage in hard regions and over-coverage in easy ones. To improve the conditional coverage, Plassier et al. [2025a] propose RCP, which transforms the raw nonconformity scores into a rectified version that is approximately homoscedastic. While the RCP framework allows for general transformations, we focus on the fundamental additive correction in our work.
@@
Since $\pi(x)$ is a function of $x$, a natural approach is to construct a metric that summarizes its deviation from the target level $1-\alpha$. Following Kiyani et al. [2024], we define the Mean Squared Conditional Error (MSCE) as:
@@
Our framework relies on estimating the conditional quantiles of the nonconformity scores. Formally, let $Z\in\mathbb{R}$ be a target random variable (typically the score $S$ in our context) and $X\in\mathcal{X}$ be the covariates. We denote the true conditional $\tau$-th quantile of $Z$ given $X=x$ as $q_{\tau}(x)\coloneqq\inf\{z:\mathbb{P}(Z\leq z\mid X=x)\geq\tau\}$ for a level $\tau\in(0,1)$. Standard quantile regression estimates $q_{\tau}$ by minimizing the expected pinball loss Koenker [2005]. The pinball loss function $\rho_{\tau}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ is defined as:
@@
In practice, given a hypothesis class $\mathcal{G}$ (e.g., neural networks), we seek an estimator $\hat{g}\in\mathcal{G}$ that minimizes the empirical risk $\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(z_{i},g(x_{i}))$. A fundamental property of this objective is that, assuming sufficient model capacity, the population minimizer $g^{*}\coloneqq\arg\min_{g}\mathbb{E}[\rho_{\tau}(Z,g(X))]$ uniquely recovers the true conditional quantile $q_{\tau}$.
@@ 3 Approximation to MSCE
In general, without a generative model, we typically observe only a single realization of $Y$ for each instance $X$. As a result, although the MSCE is a theoretically well-motivated metric, it is inherently difficult to evaluate—and hence to optimize—in practice. To address this challenge, Kiyani et al. [2024] derived an upper bound on the MSCE, while Plassier et al. [2025a] established a stronger pointwise bound on the deviation $|F_{S\mid X}(\hat{q}_{1-\alpha}(x))-(1-\alpha)|$. Though there is a gap between $F_{S\mid X}(\hat{q}_{\tau}(x))$ and $\pi(x)$ in RCP due to the conformalization step after quantile regression, Plassier et al. [2025a] proves that it shrinks at an exponential rate w.r.t. the size of sample used for conformalization.
Hereafter, we will use $\tau\coloneqq 1-\alpha$ to denote the target coverage level, and $\hat{q}_{\tau}(x)$ to denote the estimated conditional $\tau$-quantile of the nonconformity scores. We now introduce the pointwise results that motivate quantile regression with standard pinball loss Plassier et al. [2025a].
@@
Let $G(u):=\left(F_{S\mid X}(u)-\tau\right)^{2}$, and let $\mathcal{E}(x)\coloneqq\mathcal{L}_{x}(\hat{q}_{\tau}(x))-\mathcal{L}_{x}(q_{\tau}(x))$ denotes the pointwise excess risk with pinball loss. We have following two Taylor expansions with Lagrange remainder:
@@
$\mathcal{E}(x)=f_{S|X}(q(x))\epsilon_{q}(x)^{2}+\frac{1}{6}f_{S\mid X}^{\prime}\left(\xi_{S,2}\right)\epsilon_{q}(x)^{3},$ (12)
@@
$S=\nu(x)+\sigma(x)\xi,$ (15)
where $\xi$ is a standardized random variable with a base PDF $f_{0}(\cdot)$ (e.g., standard normal or Laplace) and CDF $F_{0}(\cdot)$. The conditional density of $S$ is given by $f_{S|X}(s)=\frac{1}{\sigma(x)}f_{0}\left(\frac{s-\nu(x)}{\sigma(x)}\right)$. Let $z_{\tau}=F_{0}^{-1}(\tau)$ be the $\tau$-quantile of the base distribution. The true conditional quantile of the score is then $q_{\tau}(x)=\nu(x)+\sigma(x)z_{\tau}$. Evaluating the conditional density at the true quantile yields:

### v4

@@ Abstract
While conformal prediction provides robust marginal coverage guarantees, achieving reliable conditional coverage for specific inputs remains challenging. Although exact distribution-free conditional coverage is impossible with finite samples, recent work has focused on improving the conditional coverage of standard conformal procedures. Distinct from approaches that target relaxed notions of conditional coverage, we directly minimize the mean squared error of conditional coverage by refining the quantile regression components that underpin many conformal methods. Leveraging a Taylor expansion, we derive a sharp surrogate objective for quantile regression: a density-weighted pinball loss, where the weights are given by the conditional density of the conformity score evaluated at the true quantile. We propose a three-headed quantile network that estimates these weights via finite differences using auxiliary quantile levels at $1-\alpha\pm\delta$, subsequently fine-tuning the central quantile by optimizing the weighted loss. We provide a theoretical analysis with exact non-asymptotic guarantees characterizing the resulting excess risk. Extensive experiments [^1] The code is available at https://github.com/Cqyiiii/Colorful-Pinball-Conformal-Prediction-CPCP. on diverse high-dimensional real-world datasets demonstrate remarkable improvements in conditional coverage performance.
@@ 1 Introduction
As machine learning systems are increasingly deployed in high-stakes environments, the demand for reliability extends beyond predictive accuracy to rigorous uncertainty quantification (UQ). In safety-critical domains, a model’s ability to provide appropriate predictive intervals is often as valuable as accurate point predictions. Among existing UQ frameworks, Conformal Prediction (CP; Vovk et al., 2005; Shafer and Vovk, 2008) has emerged as a paradigm of choice due to its mathematically grounded guarantees. Unlike Bayesian methods or ensemble techniques that often rely on strong distributional assumptions or heavy computational overhead, CP offers a distribution-free, model-agnostic framework. It constructs prediction sets or intervals that provably contain the ground truth with a user-specified probability $1-\alpha$ (e.g., 90%) in finite samples, providing a layer of statistical trust that is indispensable for real-world deployment.
However, standard split conformal prediction only provides a marginal guarantee over the population and cannot guarantee conditional coverage for specific instances—precisely what practitioners require in high-stakes scenarios. Although hardness results for exact distribution-free conditional guarantees with finite samples are well-established (Vovk, 2012; Lei and Wasserman, 2014; Foygel Barber et al., 2021), a growing body of works targets improving the conditional coverage of conformal procedures. To this end, many works explore to relax the conditional requirement (Angelopoulos et al., 2024), e.g., group-conditional coverage (Jung et al., 2023; Ding et al., 2023).
Instead, this paper targets controlling the exact conditional coverage rather than relaxed forms. A natural objective is the MSE of conditional coverage—termed Mean Squared Conditional Error (MSCE) in Kiyani et al. (2024). We consider this metric since minimizing the MSCE provides high-probability control over conditional coverage deviations via Bernstein’s inequality. We build upon the connection between MSCE and the excess risk of the pinball loss established in recent literature (Kiyani et al., 2024; Plassier et al., 2025a). Specifically, these works show that the MSCE is upper-bounded by the excess risk of a quantile estimator with pinball loss, suggesting that improving the quantile estimator directly translates to better conditional coverage.
@@ 1.1 Related Works
We focus our discussion on the conditional coverage of conformal prediction. In the literature, there exists another branch of work concerning the training-conditional guarantee (Park et al., 2020; Bian and Barber, 2023; Duchi, 2025), which studies the coverage probability conditional on the specific calibration set observed. To clarify, we are specifically concerned with the coverage guarantee conditional on the covariate of the test sample, $X_{\text{test}}$.
@@
Group-conditional coverage serves as a natural approximation to the exact conditional guarantee. Here, we specifically discuss the case where groups are formed by partitioning the covariate space. Classic conformal procedures operate with pre-defined groups (Jung et al., 2023), where $h$ is a fixed discrete partition function. Beyond fixed groups, Kiyani et al. (2024) propose co-training the partition function with quantile regression within each group, providing a discrete approximation of conditional coverage. Though straightforward, the precision of this discrete approximation is limited, and the convergence rate of the resulting MSCE scales as slowly as $O(n^{-1/4})$. Adopting a different perspective that reformulates the conditional guarantee as moment equations, Gibbs et al. (2025a) extends the approximation to infinite, overlapping groups. They leverage regularized quantile regression as well as full conformal prediction (Vovk et al., 2005). However, Gibbs et al. (2025a) provides control only on a substitute metric—coverage under covariate shift—rather than the exact conditional coverage. Moreover, full conformal procedures incur a significant computational burden, limiting their general applicability.
@@
Beyond group-conditional guarantees, localized conformal prediction (Bian and Barber, 2023; Hore and Barber, 2024) provides another avenue to approximate the conditional guarantee by applying kernel weighting during empirical quantile computation. As analyzed in Hore and Barber (2024), the guarantee provided by such localization corresponds to a random function $h$ that outputs a point sampled from a distribution centered at $X_{\text{test}}$, defined by the kernel. However, this strategy suffers inherently from the curse of dimensionality, particularly in the complex, high-dimensional scenarios that modern conformal prediction targets.
@@
Complementing the focus on grouping and quantile computation, a substantial body of literature strives to refine the nonconformity score function. Improving upon the classic residual score, Lei et al. (2018) propose scaled residuals using a learned conditional standard deviation $\sigma(x)$. Romano et al. (2019) propose Conformalized Quantile Regression (CQR), which directly conformalizes two learned quantiles rather than a mean prediction. They also highlighted the difficulty of learning $\sigma(x)$, particularly for overfitted neural networks. More recently, Xie et al. (2024) propose a boosting procedure to iteratively refine the score function. However, besides high computational costs, this method further requires access to training data, which is often infeasible given the prevalence of pre-trained black-box models. Similarly, CQR necessitates replacing the traditional conditional mean regression objective with quantile regression on the training set, thereby limiting its general applicability. While the aforementioned scores are tailored for one-dimensional labels, another branch of work investigates density-based scores (Izbicki et al., 2022; Plassier et al., 2025b; Braun et al., 2025a), which are readily extensible to multi-dimensional labels and also flexible enough to capture heteroscedasticity. However, estimating the conditional density $f(Y\mid X)$ is generally a harder task than regression on the conditional mean $\mathbb{E}[Y\mid X]$. Furthermore, the calibration set used for density estimation is typically much smaller than the training set. This sample size mismatch limits the efficacy of density estimation; therefore, we focus on the quantile regression of scores rather than density estimation.
@@
Most relevant to our work is the approach of performing quantile regression directly on conformity scores (Gibbs et al., 2025a; Kiyani et al., 2024; Plassier et al., 2025a). As aforementioned, the excess risk of the pinball loss constitutes an upper bound on the MSCE. Specifically, Plassier et al. (2025a) propose a general framework called Rectified Conformal Prediction (RCP), which transforms the original nonconformity score into a rectified version that tries to remove the covariate-dependent part in the score and thereby naturally improves conditional coverage of standard split conformal prediction with these rectified scores. However, naive quantile regression using the pinball loss exhibits inherent limitations in conditional coverage (Feldman et al., 2021), and struggles even with marginal coverage in high-dimensional settings (Gibbs et al., 2025b). In our context, we identify that naive quantile regression overlooks a critical heteroscedastic component. Our method aims to recover this missing component, yielding a significantly sharper approximation to the MSCE.
@@ 2 Preliminaries
We consider the standard setting where we observe i.i.d. data points $(X,Y)\in\mathcal{X}\times\mathcal{Y}$. In the split conformal prediction framework (Papadopoulos et al., 2002), the available data is randomly partitioned into two disjoint subsets: a training set $\mathcal{D}_{\text{train}}$ and a calibration set $\mathcal{D}_{\text{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ of size $n$. A predictive model is first fitted on $\mathcal{D}_{\text{train}}$. We then define a nonconformity score function $S:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, which measures the discrepancy between the target $y$ and the model’s prediction at $x$ (e.g., the absolute residual $S(x,y)=|y-\hat{\mu}(x)|$). The scores are computed for all calibration points as $S_{i}=S(X_{i},Y_{i})$ for all $i\in\{1,\dots,n\}$. Given a user-specified miscoverage level $\alpha\in(0,1)$, we compute the conformal threshold $\hat{q}$ as the $\lceil(n+1)(1-\alpha)\rceil/n$-th empirical quantile of the calibration scores $\{s_{1},\dots,s_{n}\}$. For a new test input $X_{n+1}$, the prediction set is constructed as $\mathcal{C}(X_{n+1})=\{y\in\mathcal{Y}:S(X_{n+1},y)\leq\hat{q}\}$. This procedure satisfies:
@@
As has been widely noted, the failure pattern of standard split conformal prediction is under-coverage in hard regions and over-coverage in easy ones. To improve the conditional coverage, Plassier et al. (2025a) propose RCP, which transforms the raw nonconformity scores into a rectified version that is approximately homoscedastic. While the RCP framework allows for general transformations, we focus on the fundamental additive correction in our work.
@@
Since $\pi(x)$ is a function of $x$, a natural approach is to construct a metric that summarizes its deviation from the target level $1-\alpha$. Following Kiyani et al. (2024), we define the Mean Squared Conditional Error (MSCE) as:
@@
Our framework relies on estimating the conditional quantiles of the nonconformity scores. Formally, let $Z\in\mathbb{R}$ be a target random variable (typically the score $S$ in our context) and $X\in\mathcal{X}$ be the covariates. We denote the true conditional $\tau$-th quantile of $Z$ given $X=x$ as $q_{\tau}(x)\coloneqq\inf\{z:\mathbb{P}(Z\leq z\mid X=x)\geq\tau\}$ for a level $\tau\in(0,1)$. Standard quantile regression estimates $q_{\tau}$ by minimizing the expected pinball loss (Koenker, 2005). The pinball loss function $\rho_{\tau}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ is defined as:
@@
In practice, given a hypothesis class $\mathcal{G}$ (e.g., neural networks), we seek an estimator $\hat{g}\in\mathcal{G}$ that minimizes the empirical risk $\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(z_{i},g(x_{i}))$. A fundamental property of this objective is that, assuming sufficient model capacity, the population minimizer $g^{\star}\coloneqq\arg\min_{g}\mathbb{E}[\rho_{\tau}(Z,g(X))]$ uniquely recovers the true conditional quantile $q_{\tau}$.
@@ 3 Approximation to MSCE
In general, without a generative model, we typically observe only a single realization of $Y$ for each instance $X$. As a result, although the MSCE is a theoretically well-motivated metric, it is inherently difficult to evaluate—and hence to optimize—in practice. To address this challenge, Kiyani et al. (2024) derived an upper bound on the MSCE, while Plassier et al. (2025a) established a stronger pointwise bound on the deviation $|F_{S\mid X}(\hat{q}_{1-\alpha}(x))-(1-\alpha)|$. Though there is a gap between $F_{S\mid X}(\hat{q}_{\tau}(x))$ and $\pi(x)$ in RCP due to the conformalization step after quantile regression, Plassier et al. (2025a) proves that it shrinks at an exponential rate w.r.t. the size of sample used for conformalization.
Hereafter, we will use $\tau\coloneqq 1-\alpha$ to denote the target coverage level, and $\hat{q}_{\tau}(x)$ to denote the estimated conditional $\tau$-quantile of the nonconformity scores. We now introduce the pointwise results that motivate quantile regression with standard pinball loss (Plassier et al., 2025a).
@@
Let $G(u):=\left(F_{S\mid X}(u)-\tau\right)^{2}$, and let $\mathcal{E}(x)\coloneqq\mathcal{L}_{x}(\hat{q}_{\tau}(x))-\mathcal{L}_{x}(q_{\tau}(x))$ denote the pointwise excess risk under pinball loss. Under standard regularity assumpions, we derive two Taylor expansions:
@@
$\mathcal{E}(x)=\frac{1}{2}f_{S|X}(q_{\tau}(x))\epsilon_{q}(x)^{2}+\frac{1}{6}f_{S\mid X}^{\prime}(\xi_{S,2})\epsilon_{q}(x)^{3},$ (12)
@@
$S=\varpi(x)+\sigma(x)\xi,$ (15)
where $\xi$ is a standardized random variable with a base PDF $f_{0}(\cdot)$ (e.g., standard normal or Laplace) and CDF $F_{0}(\cdot)$. The conditional density of $S$ is given by $f_{S|X}(s)=\frac{1}{\sigma(x)}f_{0}\left(\frac{s-\varpi(x)}{\sigma(x)}\right)$. Let $z_{\tau}=F_{0}^{-1}(\tau)$ be the $\tau$-quantile of the base distribution. The true conditional quantile of the score is then $q_{\tau}(x)=\varpi(x)+\sigma(x)z_{\tau}$. Evaluating the conditional density at the true quantile yields: